OpenStudy (anonymous):

Why does the derivative of f(x)=sinxcosx equal f'(x)=cos2x and not f'(x)=cos^2x-sin^2x ?

OpenStudy (anonymous):

By identity rule cos^2x-sin^2x = cos2x

OpenStudy (raden):

yea, that is similar

OpenStudy (anonymous):

Wait, but isn't sin^2x+cos^x=1, sin^2x=1-cos^2x, and cos^2x=1-sin^2x? I don't see how cos^2x-sin^2x= cos2x

OpenStudy (raden):

use this formula : cos(a+b)=cos(a)cos(b)-sin(a)sin(b) so, cos(2x)=cos(x+x)=cos(x)cos(x)-sin(x)sin(x)=(cos(x))^2-((sin(x))^2

OpenStudy (anonymous):

Whoa, awesome! Thanks! :D

OpenStudy (raden):

welcome :D